OpenStudy (anonymous):

How would you get this into Vertex Form: y+60x+294=-3x^2?

OpenStudy (saifoo.khan):

y = -3x^2 - 60x - 294. Can u solve the RIGHT side now?

OpenStudy (anonymous):

Yes, but I got 3x^2 +60x+294=-y and I don't know what to do after this.

OpenStudy (saifoo.khan):

No, you can't take y as negative.

OpenStudy (saifoo.khan):

Have a look, http://www.youtube.com/watch?v=TV5kDqiJ1Os

OpenStudy (anonymous):

This is what I did: \[y+60x+294=-3x^2\] and I took away y from each side and got:\[60x+294=-3x^2-y\]

OpenStudy (saifoo.khan):

y = ax^2 + bx + c U have to bring to this form^

OpenStudy (anonymous):

then I added 3x^2 to both sides and got: 3x^2 +60x+294= -y.

OpenStudy (saifoo.khan):

now, multiply the whole equation by -1, so that u can get y as positive.

OpenStudy (anonymous):

So it would be this: -3x^2 -60x-294=y?

OpenStudy (saifoo.khan):

YES! correct.

OpenStudy (saifoo.khan):

Now solve the equation

OpenStudy (anonymous):

Okay. Thanks for your help.

OpenStudy (saifoo.khan):

No problem